OpenStudy (anonymous):

The value of one trigonometric function is give 0

OpenStudy (anonymous):

Okay, so here you can use your identities.\[\sin^2(x) + \cos^2(x) = 1,\]\[\sin(t) = \sqrt{1-\cos^2(t),}\] and from here, you can conclude tan(t)\[\tan(t) = \sin(t)/\cos(t),\]\[\csc(t) = 1/\sin(t)\] and finally, cot(t):\[\cot(t) = 1/\tan(t).\] Plug the values you have into these equations and you'll get the values for each value of t.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

No problem!